OpenStudy (anonymous):

If a ball is thrown vertically upward from the roof of a 48 foot tall building with a velocity of 80 ft/sec, its height in feet after t seconds is s(t)=48+80t−16t2 What is the velocity of the ball when it hits the ground (height 0)?

OpenStudy (anonymous):

Do you know derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Chlorophyll can you help me out?

OpenStudy (anonymous):

i really hate word problems

OpenStudy (anonymous):

80-32 t

OpenStudy (anonymous):

You need to learn writing in equation: v = - 32t + 80 From the height = 0, <=> s(t) = ...... = 0 => t = ...? Then plug this t value into v!

OpenStudy (anonymous):

what do you mean <=>?

OpenStudy (anonymous):

Height = 0 -16t² + 80t + 48 = 0

OpenStudy (anonymous):

ah ok got it thanks

OpenStudy (anonymous):

im sorry can be a little slow at times :P

OpenStudy (anonymous):

I'm trying to show you how to grasp the concept, so you'll able to be confident for testing time :)

OpenStudy (anonymous):

lol midterm was today and i forgot my calculator. didnt go very well :(

OpenStudy (anonymous):

because professor is a wingspan and wouldnt let me

OpenStudy (anonymous):

he isnt nice

OpenStudy (anonymous):

ya dont have a graphing calc yet so dont need batteries, and always have it with me but left it in my room because i fell asleep while doing my hw yesterday

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

sorry but can you help me set this up? The radius of a spherical balloon is increasing at a rate of 2 centimeters per minute. How fast is the volume changing when the radius is 10 centimeters? Note: The volume of a sphere is given by V=(4/3)πr^3. i know what to do just not how to do it

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

V = (4/3) π r³ -> v' = 4 π r² r' When r' = 2cm/ min, r = 10 cm, => v' = 4 * 100 * (+2) = ....

OpenStudy (anonymous):

I meant : => v' = 4 * 100 * (+2) π = ....

OpenStudy (anonymous):

402pi?

OpenStudy (anonymous):

oh nvm 800pi

OpenStudy (anonymous):

ya just added the 2 instead of multiplying it

OpenStudy (anonymous):

the next question could i try going through it step by step with you to make sure i have it?

OpenStudy (anonymous):

No, multiply with ( positive 2 since the rate Increasing) -> v' = 4 π r² r' => v' = 4 * 100 * (+2) π

OpenStudy (anonymous):

Okie, but open the new post, though!

OpenStudy (anonymous):

is it just me or is this site really glitchie?

OpenStudy (anonymous):

No, that's the OS site bug! My screen sometime jumping sometime froze out of control :(

OpenStudy (anonymous):

ya ok brb